OpenStudy (anonymous):

(6.1, 3.9) and (4.2, 8.1) Estimate the midpoint of the points

OpenStudy (hoblos):

x= (4.2+6.1)/2 = 10.3/2 = 5.15 y=(3.9+8.1)/2 = 12/2=6

OpenStudy (hoblos):

so the midpoint is (5.15,6)

OpenStudy (apoorvk):

for any two points (x,y) and (x',y'), their midpoints are, [(x+x')/2 , (y+y')/2] so ans is (5.15,6)

OpenStudy (anonymous):

thanz

OpenStudy (hoblos):

yw =)

OpenStudy (anonymous):

wat do yw

OpenStudy (hoblos):

yw = you are welcome :D:D